OpenStudy (anonymous):

What is the value of the expression when a = 2? 4 5a

OpenStudy (anonymous):

@pooja195

OpenStudy (moazzam07):

Please right the question neatly

OpenStudy (anonymous):

? that is the whole thing

OpenStudy (steve816):

5(2)=10

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@GracieBugg

Nnesha (nnesha):

\[\huge\rm \frac{4 }{ 5a }\] is this right ?

OpenStudy (anonymous):

yes well it is not a fraction it is 5a and 4 on top

OpenStudy (anonymous):

like 4 top side

Nnesha (nnesha):

if it's not a fraction then top of what ?

Nnesha (nnesha):

ohhh \[\huge\rm 5a^4\] like this

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

like that

Nnesha (nnesha):

so use ^<-- this mark for the exponent

Nnesha (nnesha):

now substitute a for 2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and it is 2?

OpenStudy (anonymous):

how do u do that?

OpenStudy (anonymous):

i do not get this i need help what do i do?

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

just replace a with 2\[5(2)^4\]

OpenStudy (anonymous):

ok and what next

OpenStudy (anonymous):

ad?

OpenStudy (anonymous):

add

Nnesha (nnesha):

(2)^4 is same as 2 times 2 times 2 times 2 so multiply them

OpenStudy (anonymous):

so 2*4?

OpenStudy (anonymous):

cut to the chase

OpenStudy (anonymous):

hello?

Nnesha (nnesha):

2^4 is same as \[2 \times 2 \times 2 \times 2 \]

Nnesha (nnesha):

multiply them1!!!!

OpenStudy (anonymous):

8

Nnesha (nnesha):

u sure ?

OpenStudy (anonymous):

idk i did 2*2*2*2

OpenStudy (anonymous):

and got 8

Nnesha (nnesha):

that's not right i have no idea how u r multiplying them.

OpenStudy (anonymous):

just give me the answer i will meadle u

Nnesha (nnesha):

ohh really will u give me the medal ??

OpenStudy (anonymous):

yep

Nnesha (nnesha):

you really did Oh my god what would i do with it ? eh ?

Nnesha (nnesha):

now multiply them

Nnesha (nnesha):

2^4 means 4 2's \[\huge\rm 2 \times 2 \times 2 \times 2=?\]

OpenStudy (anonymous):

16?

Nnesha (nnesha):

yes right \[5(16) = ?\]

OpenStudy (anonymous):

5(16)2?

Nnesha (nnesha):

how did you get 2 ?

OpenStudy (anonymous):

idk guesstd it

Nnesha (nnesha):

\[\huge\rm 5(\color{Red}{2)^4}\]\[\huge\rm 5(\color{Red}{16)}\] multiply them

OpenStudy (anonymous):

32

Nnesha (nnesha):

yes!!

OpenStudy (anonymous):

32 is my answer

Nnesha (nnesha):

better to use calculator instead asking for direct answer. why waste time ?

OpenStudy (anonymous):

ok thx thow

Nnesha (nnesha):

np :)

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @shadybubba 32 is my answer \(\color{blue}{\text{End of Quote}}\) no

Nnesha (nnesha):

wait what 5(16) = ??

OpenStudy (anonymous):

rong it was 80!!!!!!!

Nnesha (nnesha):

yes

OpenStudy (anonymous):

! the answer waqs 80 ur just going around telling ppl worg answers

OpenStudy (anonymous):

u said it was 32

Nnesha (nnesha):

you re the one who i said 32

OpenStudy (anonymous):

ur supost to correct me

OpenStudy (anonymous):

so ima report u it sayes in the code of conduct not to go around telling people wrong answers sorry i need to thx for wasting my time

Nnesha (nnesha):

good for you.

OpenStudy (anonymous):

thx

Nnesha (nnesha):

and it's h.w question so it's okay

OpenStudy (anonymous):

nope sorry

Nnesha (nnesha):

wat nope ? it's a test or quiz ?

Nnesha (nnesha):

alright i apologize :)

OpenStudy (freckles):

It isn't @Nnesha 's fault. The assignment is yours and it your responsibility to double check @shadybubba . No one is responsible for your work but you. @Nnesha gave get great assistance and you just made guesses such as the one to 2*2*2*2 ... You said 8... He even asked you to compute 5(16) which would have given you the correct result if you have tried to actually compute 5(16)...